The length of time patients must wait to see a doctor at an emergency room of a large hospital is uniformly distributed between 40 minutes and 3 hours. What is the probability that a patient will have to wait between 50 minutes and 1.5 hours to see a doctor?



1
Expert's answer
2022-04-21T11:42:20-0400

Note that 3 hours = 180 minutes and 1.5 hours = 90 minutes.

Since the distribution is uniform, we can find the probability by dividing the length of given period:

90 min - 50 min = 40 min

by the total length of waitng:

180 min - 40 min = 140 min,


P = 40/140 = 2/7 = 0.2857